# SFormer-OR: Single-Stage Bimodal Transformer for Scene Graph Generation in OR

### v1

@@ 3.3 Relation-Sensitive Transformer
Unlike previous works [31, 6] that predict relationships via GCN, we input dynamic relation queries into the relation-sensitive transformer block to predict predicates between subject-object pairs. As illustrated in Fig. 3, the dynamic relational query is first fed into the self-attention module to enhance the relational representation of each triplet, obtaining intermediate features . Considering that relation queries are insensitive to location information during self-attention operation, we do not input position encoding [29] into the transformer block. After that, we utilize cross-attention operation to interact with 2D semantic features. Here, we use as the Query and the unified feature is regarded as the encoder memory feature for the Key and Value. Since (View#4) concentration on the operation relationship between the surgeon and the patient, we also incorporate (View#4) features into the memory features to optimize predictions of some action relationships (e.g., touching, suturing, and cementing). As shown in Fig. 2, we concatenate the View#4 features extracted from the last layer of ResNet with the unified feature: =. Combining CNN features essentially serves to supplement encoder memory features with lower-level features. The memory feature is used as the Key and Value. Sec. 4.3 discusses the effectiveness of this operation in detail with the ablation study. Additionally, to improve the efficiency of cross-attention, inspired by masked attention from Mask2Former [32], we retain only the union of foreground regions of entities for the attention operation. The relation cross-attention operation can be expressed as
@@ 3.4 Objective Function
We train our model in an end-to-end manner. The total objective function of SFormer-OR consists of the entity box loss and the relation classification loss in the relationship-sensitive decoder:
where is the weighting factor. In line with DETR, contains L1 loss and GIoU loss [33] for box regression and Focal loss [34] () for entity classification. Here, is set to 1. In addition, we add the auxiliary loss [35] after each decoder layer. The relational classification loss is mainly implemented by Focal loss:
@@ 4.1.2 Evaluation Metrics
Consistent with the state-of-the-art methods [6, 5], we used precision, recall, and Macro F1 metrics to evaluate predicted scene graph relationships. We also employed the same evaluation metrics as [6, 5] in the downstream task of clinical role prediction. Moreover, in comparison with the generic SGG models on the 4D-OR benchmark, we adopted several widely used metrics for generic SGG model evaluation, including recall metric under the IoU threshold of 0.5 (R50), weighted mean average precision of relationship detection (wmAP), and phrase detection (wmAP) from Open Images Challenge [36] for a more comprehensive evaluation.
@@ 4.2 Implementation Details
Stage | Methods | Params | Metrics | Assist | Cement | Clean | CloseTo | Cut | Drill | Hammer | Hold | LyingOn | Operate | Prepare | Saw | Suture | Touch | Avg |
Multi-Stage | 4D-OR [6] | 84.8M | Precision | 0.42 | 0.78 | 0.53 | 0.97 | 0.49 | 0.87 | 0.71 | 0.55 | 1.00 | 0.55 | 0.62 | 0.69 | 0.60 | 0.41 | 0.68 |
Recall | 0.93 | 0.78 | 0.63 | 0.89 | 0.49 | 1.00 | 0.89 | 0.95 | 0.99 | 0.99 | 0.91 | 0.91 | 1.00 | 0.69 | 0.87 | |||
Macro F1 | 0.58 | 0.78 | 0.57 | 0.93 | 0.49 | 0.93 | 0.79 | 0.70 | 0.99 | 0.71 | 0.74 | 0.79 | 0.75 | 0.51 | 0.75 | |||
Multi-Stage | LABRAD-OR [5] | 85.8M | Precision | 0.60 | 0.96 | 0.86 | 0.96 | 0.91 | 1.00 | 0.93 | 0.71 | 1.00 | 0.85 | 0.77 | 0.78 | 1.00 | 0.71 | 0.87 |
Recall | 0.86 | 0.93 | 0.72 | 0.94 | 0.68 | 0.94 | 0.95 | 0.95 | 1.00 | 0.99 | 0.91 | 0.93 | 1.00 | 0.72 | 0.90 | |||
Macro F1 | 0.71 | 0.94 | 0.78 | 0.95 | 0.78 | 0.97 | 0.94 | 0.81 | 1.00 | 0.91 | 0.84 | 0.85 | 1.00 | 0.71 | 0.88 | |||
Single-Stage | SFormer-OR | 60.6M | Precision | 0.65 | 1.00 | 0.91 | 0.97 | 0.82 | 0.98 | 0.94 | 0.79 | 1.00 | 0.90 | 0.83 | 0.93 | 0.94 | 0.81 | 0.90 |
Recall | 0.66 | 0.94 | 0.86 | 0.93 | 0.85 | 0.92 | 0.94 | 0.83 | 1.00 | 0.91 | 0.89 | 0.99 | 0.97 | 0.49 | 0.88 | |||
@@ 4.3 Comparison with State-of-the-Arts
To evaluate the performance of our SFormer-OR for the OR-SGG task, we carried out detailed comparisons with existing SGG approaches tailored to OR scenes and generic single-stage 2D SGG models on the 4D-OR benchmark. For a fair comparison, we uniformly use the official code to train and evaluate each model on 4D-OR validation and test sets.
@@ 4.3.1 Comparison with OR-SGG Methods
We conducted a comprehensive comparison between the proposed SFormer-OR and existing OR-SGG models, i.e., 4D-OR [6] and LABRAD-OR [5]. Table 1 exhibits quantitative results of our model and competitors in terms of the precision, recall, and macro F1 metrics. We can observe that SFormer-OR demonstrates promising performance, particularly in average precision and macro F1 score, achieving 0.90 and 0.89, respectively. Benefiting from the thorough fusion of semantic features and structural information, our model excelled in predicting critical surgeon-patient actions, e.g., achieving Macro F1 scores of 0.97, 0.91, and 0.96 in cementing, operating, and sawing relations, respectively. More importantly, SFormer-OR acts as an end-to-end model with fewer parameters, enabling the direct generation of scene graphs in practical applications by feeding surgical data without intermediate prediction processes (e.g., no need for pose estimation labels). As a single-stage architecture, the inference time of our model is also faster than the 4D-OR model by 0.3s per frame, significantly increasing the potential of model deployment in real clinical practice.
To further intuitively evaluate the effectiveness of generated scene graphs, we visualize the prediction results of some keyframes in Fig. 4. Notably, we omitted the “CloseTo” relationship to explicitly display the scene graph that is most relevant to the surgical procedure. We can see that our model consistently predicts the interactions of entities without missing key messages in complex operating room environments. Compared to state-of-the-art SGG-OR models, our model exhibits more accurate detection of entities, avoiding the influence of erroneous entities in predicting relationships. While still producing redundant predictions in some cases, our model can predict the overall structure of the scene graph more precisely and robustly compared with others.
@@ 4.3.2 Comparison with Single-Stage SGG Methods
To further validate the robustness of our method, we also compare with the cutting-edge single-stage SGG methods in a generic scenario of natural scene generation (i.e., SGTR [12] and RelTR [13]). We reimplemented them on the 4D-OR dataset. Both of them are novel single-modal end-to-end architectures based on transformers. As shown in Table 2, our model not only achieves better results in vanilla metrics but also outperforms other competitors in R50, wmAP, and wmAP metrics. This can be attributed to the interaction of multi-view features via the proposed VST scheme and the extraction of geometric information from point cloud modalities by our GVC operation. Further, the sufficient utilization of relational trait priors in generating dynamic relation queries and the efficient relation-sensitive transformer structure have also brought performance gains to our model.
@@ 4.4.1 Effectiveness of Each Component
To validate the effectiveness of each component in our model, we conduct detailed ablation experiments to optimize model performance. The main components include View-Sync Transfusion (VST), Geometry-Visual Cohesion (GVC), and Relation-sensitive Transformer (RST). Note that when ablating the VST module, we replace the input of multi-view images with only the main-view image. When removing the GVC module, we directly project the extracted point cloud features to synergic features through the intrinsic and extrinsic matrices provided in the dataset. For our RST part, we adopt a GCN structure from the previous OR-SGG model [6] as a replacement to perform relation predictions. The results in Table 3 indicate that each proposed component imposes a positive impact on SFormer-OR. Compared to the baseline, our model gains 0.19, 0.24, and 0.22 improvement on the precision, recall, and macro F1 metrics after adding three essential parts (refer to the first and sixth rows). The third row of Table 3 illustrates that the GVC operation improves the generated accuracy by sufficiently integrating point cloud features into 2D semantic features. Moreover, the fifth row highlights the indispensability of RST, where the dynamic relation queries and relation-aware cross-attention operations make our model more conducive to inferring relations. Compared to other variants, SFormer-OR has no substantial increase in parameters after embedding all components, while achieving better performance.
@@ 4.4.2 Number of Relation Queries
Relation queries are an essential component in our relation-sensitive transformer. To examine the impact of the number of relation queries on relation prediction, we compare the performance of our dynamic relation queries (i.e., ) with other fixed number queries (i.e., 20, 50, 100, and 300). Unlike dynamic queries, to produce a fixed number of queries, we directly pair the entity proposals output in the entity unit without using category scores for suppression. Then, we apply an MLP operation to all entity pairs to obtain corresponding confidence values and use the top-K selection to filter out a fixed number of pairs for relation-aware query generation. As shown in Table 4, with a fixed number of relation queries, the performance of SFormer-OR constantly increases with the number of queries until the number reaches 100. Compared to the fixed-amount setting, our dynamic relation queries provide greater adaptability to variable scene interactions, making our relation-sensitive transformer more accurate in graph prediction, e.g., 0.87 vs. 0.85 in macro F1.
@@ 4.4.3 Effect of Relational Trait Priors
In the relation-sensitive decoder, we inject the latent entity pairs with various subject-object trait priors to accurately express relation queries, including semantic, spatial, and point cloud features. To validate the effectiveness of each feature representation, we compare the contributions of different combinations of prior trait embeddings on the 4D-OR validation set. As illustrated in Table 5, in addition to the indispensable semantic features, the spatial feature representation of entity pairs remarkably improves the precision value by 3%. It indicates that spatial features can connect entity pairs with the closest correlations. Additionally, point cloud features also contribute to enhancing the structural representation of entity pairs. By aggregating all relational trait priors, SFormer-OR reaches the best performance.
@@ 4.5 Clinical Role Prediction
Assigning role labels to humans in surgical scenes facilitates further improving surgical process monitoring and optimization. Since the 4D-OR benchmark provides role labels for each entity associated with the human attribute (i.e., Patient, Head Surgeon, Assistant Surgeon, Circulating Nurse, and Anaesthetist), we perform clinical role prediction based on the generated scene graphs. For a fair comparison, we also adopt Graphormer [37] to infer the clinical role probabilities of human entities as in the 4D-OR method [6]. Since our model is a single-stage framework without intermediate results of human pose estimation, we select the predicted entities belonging to the human attribute as human nodes. Fig. 6 compares the performance of our model with 4D-OR on the downstream task of clinical role prediction. It indicates that SFormer-OR achieves strong performance in all clinical roles, especially in the ‘Anaesthetist’ and ‘Circulating Nurse’ categories, improving the macro F1 score by 14% and 5%. Since the anaesthetist is usually located in the corner of operating rooms and appears exclusively at the beginning stage of surgery, there is still room for further improvement. Overall, our model carries over the advantages of scene graph generation to clinical role prediction.

### v2

@@ 3.3 Relation-Sensitive Transformer
Unlike previous works [32, 6] that predict relationships via GCN, we input dynamic relation queries into the relation-sensitive transformer block to predict predicates between subject-object pairs. As illustrated in Fig. 3, the dynamic relational query is first fed into the self-attention module to enhance the relational representation of each triplet, obtaining intermediate features . Considering that relation queries are insensitive to location information during self-attention operation, we do not input position encoding [30] into the transformer block. After that, we utilize cross-attention operation to interact with 2D semantic features. Here, we use as the Query and the unified feature is regarded as the encoder memory feature for the Key and Value. Since (View#4) concentration on the operation relationship between the surgeon and the patient, we also incorporate (View#4) features into the memory features to optimize predictions of some action relationships (e.g., touching, suturing, and cementing). As shown in Fig. 2, we concatenate the View#4 features extracted from the last layer of ResNet with the unified feature: =. Combining CNN (Convolution Neural Network) features essentially serves to supplement encoder memory features with lower-level features. The memory feature is used as the Key and Value. Sec. 4.4 discusses the effectiveness of this operation in detail with the ablation study. Additionally, to improve the efficiency of cross-attention, inspired by masked attention from Mask2Former [33], we retain only the union of foreground regions of entities for the attention operation. The relation cross-attention operation can be expressed as
@@ 3.4 Objective Function
We train our model in an end-to-end manner. The total objective function of S2Former-OR consists of the entity box loss and the relation classification loss in the relationship-sensitive decoder:
where is the weighting factor. In line with DETR, contains L1 loss and GIoU (Generalized Intersection over Union) loss [34] for box regression and Focal loss [35] () for entity classification. Here, is set to 1. In addition, we add the auxiliary loss [36] after each decoder layer. The relational classification loss is mainly implemented by Focal loss:
@@ 4.1.2 Evaluation Metrics
Consistent with the state-of-the-art methods [6, 5], we used precision, recall, and Macro F1 metrics to evaluate predicted scene graph relationships. We also employed the same evaluation metrics as [6, 5] in the downstream task of clinical role prediction. Moreover, in comparison with the generic SGG models on the 4D-OR benchmark, we adopted several widely used metrics for generic SGG model evaluation, including recall metric under the IoU threshold of 0.5 (R50), weighted mean average precision of relationship detection (wmAPrel), and phrase detection (wmAPphr) from Open Images Challenge [37] for a more comprehensive evaluation.
@@ 4.2 Implementation Details
Stage | Methods | Params | Time/s | Metrics | Assist | Cement | Clean | CloseTo | Cut | Drill | Hammer | Hold | LyingOn | Operate | Prepare | Saw | Suture | Touch | Avg |
Multi-Stage | 4D-OR [6] | 84.8M | Precision | 0.42 | 0.78 | 0.53 | 0.97 | 0.49 | 0.87 | 0.71 | 0.55 | 1.00 | 0.55 | 0.62 | 0.69 | 0.60 | 0.41 | 0.68 | |
1.28 | Recall | 0.93 | 0.78 | 0.63 | 0.89 | 0.49 | 1.00 | 0.89 | 0.95 | 0.99 | 0.99 | 0.91 | 0.91 | 1.00 | 0.69 | 0.87 | |||
Macro F1 | 0.58 | 0.78 | 0.57 | 0.93 | 0.49 | 0.93 | 0.79 | 0.70 | 0.99 | 0.71 | 0.74 | 0.79 | 0.75 | 0.51 | 0.75 | ||||
LABRAD-OR [5] | 85.8M | Precision | 0.60 | 0.96 | 0.86 | 0.96 | 0.91 | 1.00 | 0.93 | 0.71 | 1.00 | 0.85 | 0.77 | 0.78 | 1.00 | 0.71 | 0.87 | ||
1.46 | Recall | 0.86 | 0.93 | 0.72 | 0.94 | 0.68 | 0.94 | 0.95 | 0.95 | 1.00 | 0.99 | 0.91 | 0.93 | 1.00 | 0.72 | 0.90 | |||
Macro F1 | 0.71 | 0.94 | 0.78 | 0.95 | 0.78 | 0.97 | 0.94 | 0.81 | 1.00 | 0.91 | 0.84 | 0.85 | 1.00 | 0.71 | 0.88 | ||||
Single-Stage | S2Former-OR | 60.6M | Precision | 0.65 | 1.00 | 0.91 | 0.97 | 0.82 | 0.98 | 0.94 | 0.79 | 1.00 | 0.90 | 0.83 | 0.93 | 0.94 | 0.81 | 0.90 | |
0.98 | Recall | 0.66 | 0.94 | 0.86 | 0.93 | 0.85 | 0.92 | 0.94 | 0.83 | 1.00 | 0.91 | 0.89 | 0.99 | 0.97 | 0.49 | 0.88 | |||
@@ 4.3 Comparison with State-of-the-Arts
To evaluate the performance of our S2Former-OR for the OR-SGG task, we carried out detailed comparisons with existing SGG approaches tailored to OR scenes and generic single-stage 2D SGG models on the 4D-OR benchmark. For a fair comparison, we uniformly use the official code to train and evaluate each model on 4D-OR validation and test sets.
@@ 4.3.1 Comparison with OR-SGG Methods
We conducted a comprehensive comparison between the proposed S2Former-OR and existing OR-SGG models, i.e., 4D-OR [6] and LABRAD-OR [5]. Table 1 exhibits quantitative results of our model and competitors in terms of the precision, recall, and macro F1 metrics. We can observe that S2Former-OR demonstrates promising performance, particularly in average precision and macro F1 score, achieving 0.90 and 0.89, respectively. Benefiting from the thorough fusion of semantic features and structural information, our model excelled in predicting critical surgeon-patient actions, e.g., achieving Macro F1 scores of 0.97, 0.91, and 0.96 in cementing, operating, and sawing relations, respectively. More importantly, the proposed S2Former-OR, acting as a bi-modal end-to-end model, is capable of directly generating scene graphs by feeding surgical data without intermediate prediction processes (e.g., no need for pose estimation). Furthermore, As illustrated in the third and fourth columns of Table 1, our single-stage model has advantages over multi-stage architectures in terms of parameters and inference time, e.g., 0.3 seconds per frame faster than the 4D-OR model, significantly boosting the potential of model deployment in real clinical practice.
To further intuitively evaluate the effectiveness of generated scene graphs, we visualize the prediction results of some keyframes in Fig. 4. Notably, we omitted the “CloseTo” relationship to explicitly display the scene graph that is most relevant to the surgical procedure. We can see that our model consistently predicts the interactions of entities without missing key messages in complex operating room environments. Compared to state-of-the-art OR-SGG models, our model exhibits more accurate detection of entities, avoiding the influence of erroneous entities in predicting relationships. While still producing redundant predictions in some cases, our model can predict the overall structure of the scene graph more precisely and robustly compared with others.
@@ 4.3.2 Comparison with Single-Stage SGG Methods
To further validate the robustness of our method, we also compare with the cutting-edge single-stage SGG methods in a generic scenario of natural scene generation, including SGTR [12], RelTR [13], and Structured Sparse R-CNN [26]. We retrain and validate these novel single-modal end-to-end architectures on the 4D-OR dataset. For a fair comparison, we ablate the values of learning rate and batch size for each model, and other components remain the same as the default settings. As shown in Table 2, our model not only achieves better results in vanilla metrics but also outperforms other competitors in R50, wmAPrel, and wmAPphr metrics. This can be attributed to the interaction of multi-view features via the proposed VST scheme and the extraction of geometric information from point cloud modalities by our GVC operation. Further, the sufficient utilization of relational trait priors in generating dynamic relation queries and the efficient relation-sensitive transformer structure have also brought performance gains to our model.
Methods | Backbones | R50 | wmAPrel | wmAPphr | Prec | Rec | F1 |
SGTR [12] | ResNet-101 | 0.57 | 0.51 | 0.69 | 0.69 | 0.70 | 0.69 |
RelTR [13] | ResNet-50 | 0.54 | 0.48 | 0.62 | 0.61 | 0.67 | 0.63 |
SS R-CNN [26] | ResNeXt-101 | 0.64 | 0.55 | 0.74 | 0.82 | 0.79 | 0.80 |
S2Former-OR | ResNet-50 | 0.71 | 0.63 | 0.83 | 0.90 | 0.84 | 0.87 |
@@ 4.4.1 Effectiveness of Each Component
To validate the effectiveness of each component in our model, we conduct detailed ablation experiments to optimize model performance. The main components include View-Sync Transfusion (VST), Geometry-Visual Cohesion (GVC), and Relation-sensitive Transformer (RST). Note that when ablating the VST module, we replace the input of multi-view images with only the main-view image. When removing the GVC module, we directly project the extracted point cloud features to synergic features through the intrinsic and extrinsic matrices provided in the dataset. For our RST part, we adopt a GCN structure from the previous OR-SGG model [6] as a replacement to perform relation predictions. The results in Table 3 indicate that each proposed component imposes a positive impact on S2Former-OR. Compared to the baseline, our model gains 0.19, 0.24, and 0.22 improvement on the precision, recall, and macro F1 metrics after adding three essential parts (refer to the first and sixth rows). The third row of Table 3 illustrates that the GVC operation improves the generated accuracy by sufficiently integrating point cloud features into 2D semantic features. Moreover, the fifth row highlights the indispensability of RST, where the dynamic relation queries and relation-aware cross-attention operations make our model more conducive to inferring relations. Compared to other variants, S2Former-OR has no substantial increase in parameters after embedding all components, while achieving better performance.
@@
Q,K,V in VST | Prec | Rec | F1 |
Q: View#2 K,V: View#1,3,4 | 0.82 | 0.77 | 0.79 |
Q: View#3 K,V: View#1,2,4 | 0.81 | 0.77 | 0.79 |
Q: View#4 K,V: View#1,2,3 | 0.83 | 0.81 | 0.82 |
Q: View#1 K,V: View#2 | 0.86 | 0.81 | 0.83 |
Q: View#1 K,V: View#2,3 | 0.87 | 0.83 | 0.85 |
Q: View#1 K,V: View#2,3,4 | 0.90 | 0.84 | 0.87 |
@@ 4.4.2 Number of Relation Queries
To examine the impact of the number of relation queries on relation prediction, we compare the performance of our dynamic relation queries (i.e., ) with other fixed number queries (i.e., 20, 50, 100, 150, 200, and 300). Unlike dynamic queries, to produce a fixed number of queries, we directly pair the entity proposals output in the entity unit without using category scores for suppression. Then, we apply an MLP operation to all entity pairs to obtain corresponding confidence values and use the top-K selection to filter out a fixed number of pairs for relation-aware query generation. As shown in Table 4, with a fixed number of relation queries, the performance of S2Former-OR constantly increases with the number of queries until the number reaches 150. For dynamic relation queries, we perform distribution statistics on the number of entities for each sample from 4D-OR, where the mean value of is 12.29 and the median value is 13. Thus, the number of dynamic relation queries N*(N-1) is approximately 13(13-1)=156, which is close to the optimal fixed number of relation queries at 150. Compared to the fixed-amount setting, our dynamic relation queries provide greater adaptability to variable scene interactions, making our relation-sensitive transformer more accurate in graph prediction.
4.4.3 Different combinations of viewpoints in VST
In the proposed VST scheme, we interact the main-view features with the local fine-grained information of other multiple views to enhance the holistic 2D semantic representation and mitigate occlusions. To investigate the effect of different viewpoint features on relationship prediction of S2Former-OR, we conduct an ablation study on combinations of multi-view features in the VST scheme. The experimental results in Table 5 show that multi-view information, especially View#1 and View#2, can effectively facilitate scene graph generation in ORs. In addition, the selection of the Query, Key, and Value of VST is also critical. Our model performs better when the features of View#1 are served as Queries and the features from other views are used as Keys and Values. This can be attributed to the main view (View#1) providing global features that contain various entities and relationships, which is more suitable for feature transfusion as the Query. Other views focus more on specific locations and relationship classes, making them more appropriate as Keys and Values to complement the features that round out the main view.
@@ 4.4.4 Effect of Relational Trait Priors
In the relation-sensitive decoder, we inject the latent entity pairs with various subject-object trait priors to accurately express relation queries, including semantic, spatial, and point cloud features. To validate the effectiveness of each feature representation, we compare the contributions of different combinations of prior trait embeddings on the 4D-OR validation set. As illustrated in Table 6, in addition to the indispensable semantic features, the spatial feature representation of entity pairs remarkably improves the precision value by 3 percentage points. It indicates that spatial features can connect entity pairs with the closest correlations. Additionally, point cloud features also contribute to enhancing the structural representation of entity pairs. By aggregating all relational trait priors, S2Former-OR reaches the best performance.
@@ 4.4.5 Memory Feature in Relation-Sensitive Transformer
Number of Points | Prec | Rec | F1 | Time/s |
256 | 0.88 | 0.75 | 0.81 | 0.95 |
512 | 0.89 | 0.84 | 0.86 | 0.96 |
1024 | 0.90 | 0.84 | 0.87 | 0.98 |
2048 | 0.89 | 0.85 | 0.87 | 1.02 |
4.4.6 Number of Point Cloud Features
Downsampling to different numbers of point cloud features using PointNet++ [27] is crucial for subsequent geometry-visual cohesion (GVC) operation. To this end, we conduct an ablation experiment in Table 8 for different numbers of point cloud features input to the GVC operation, including 256, 512, 1024, and 2048 sampling points. The experimental results demonstrate that as the number of sampling points increases, the contribution of 3D geometric information to the accuracy of scene graph prediction also gradually improves. The performance of our model reaches a plateau when the number of sampling points reaches or exceeds 1024. In order to achieve a better balance between accuracy and computational efficiency, we finally select 1024 points as the default point cloud features for our S2Former-OR.
@@ 4.5 Clinical Role Prediction
Assigning role labels to humans in surgical scenes facilitates further improving surgical process monitoring and optimization. Since the 4D-OR benchmark provides role labels for each entity associated with the human attribute (i.e., Patient, Head Surgeon, Assistant Surgeon, Circulating Nurse, and Anaesthetist), we perform clinical role prediction based on the generated scene graphs. For a fair comparison, we also adopt Graphormer [38] to infer the clinical role probabilities of human entities as in the 4D-OR method [6]. Since our model is a single-stage framework without intermediate results of human pose estimation, we select the predicted entities belonging to the human attribute as human nodes. Fig. 6 compares the performance of our model with 4D-OR on the downstream task of clinical role prediction. It indicates that S2Former-OR achieves strong performance in all clinical roles, especially in the ‘Anaesthetist’ and ‘Circulating Nurse’ categories, improving the macro F1 score by 14 and 5 percentage points, respectively. Since the anaesthetist is usually located in the corner of operating rooms and appears exclusively at the beginning stage of surgery, there is still room for further improvement. Overall, our model carries over the advantages of scene graph generation to clinical role prediction.
4.6 Performance on 3DSSG Benchmark
To demonstrate the generalization capabilities of S2Former-OR and to extend the impact of this work, we also verify the performance of our model on the 3D semantic scene graph (3DSSG) benchmark [39] for the 3D indoor scene graph generation. The dataset provides 3D point cloud samples, multi-view RGB inputs, and semantic scene graph annotations. In this experiment, we use the same data preparation and training/validation partitioning as 3DSSG. The labeling of the 3D instance is known based on the setup of existing related methods [39, 15] in 3DSSG. Thus, we can also project the 3D instance labels into the 2D space by using the camera parameters, thereby obtaining 2D object locations. Then, we directly utilize the position of instances instead of the entity prediction head to generate latent entity pairs. Other settings remain unchanged. We compare our method with several well-known 3D SGG methods on the 3DSSG validation set, including SGPN [39] SGFN [40], and VL-SAT [15]. For the evaluation metrics, we use the top-k accuracy (A50, A100) and the mean top-k accuracy (mAk) to evaluate the triplet predictions. As shown in Table 9, our model also achieves competitive results on the indoor 3DSSG dataset, especially on the mA50 and mA100 metrics, demonstrating the strong generalization capability of our S2Former-OR across general multimodal SGG tasks.